OpenStudy (anonymous):

write each trigonometric expression in terms of a single trignometric function. 1-2(sin^2)5ß

OpenStudy (paxpolaris):

\[\cos(2\theta)=\cos^2\theta-\sin^2\theta\\ \ \ \ \ \ \ \ \ \ \ \ \ =1-2\sin^2\theta\]

OpenStudy (paxpolaris):

use this formula in reverse for: \[1-2\sin^2\left( 5\beta \right)\]